Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Why is it always 1-(Probability of it NOT Occurring)?

Expert replies
by GTwinkie » Sat Dec 06, 2008 9:56 pm
A machine consists of three components, which are components X, Y, and Z. These three components operate properly independently of one another. The machine operates properly only when component X operates properly and at least one of components Y and Z operates properly. The probability that component X operates properly is 0.8, the probability that component Y operates properly is 0.4, and the probability that component Z operates properly is 0.3. What is the probability that the machine operates properly?

A) 0.096
B) 0.252
C) 0.464
D) 0.583
E) 0.648

In all the Kaplan explanations for probability questions, it always says, "It is easier for us to find the probability of XYZ not occurring".

Can anyone explain why that is? Why can't we just find the probability of XYZ actually occurring?

For example, in the question above, they tell us outwardly the probability of Components Y and Z operating properly, 0.4 and 0.3 respectively. My inclination is to simply multiply the probability of Component X working (0.8) with the probabilities of Y and Z working, 0.8 * 0.4 * 0.3 = 0.096. While, I know that sounds unreasonably low and is ultimately the incorrect answer, why doesn't this simple approach work?

The explanation states that you must first find the probability of Y and Z NOT operating properly, i.e.

1 - 0.4 = 0.6
1 - 0.3 = 0.7

Then multiply them to get 0.42 and subtract it from 1 to find the probability of either Y or Z operating properly, which is 0.58. At that point you just multiply 0.58 by 0.8 and the answer is 0.464.

If anyone can shed light on this, I'd appreciate it. Thanks everyone!
Join the discussion
Source: — Problem Solving |

by Neo2000 » Sun Dec 07, 2008 11:00 am
The question states that "X and atleast one of Y or Z"

You therefore have a number of situtations

X works,Y doesnt, Z does
X Works, Y does, Z doesnt
X works, Y works, Z works

What you are doing is considering only the third situation without considering the first 2

Since we know that total probability = 1
Finding out the sum of cases that dont work and subtracting that from 1 gives you the probability that does work
Join the discussion

by Jen Kedro » Wed Dec 10, 2008 12:41 pm
Exactly...since the probability of something happening, and the probability of that same thing NOT happening, add to 1 (i.e. if probability it will rain is 30%, probability it won't rain is 70%, or total of 100%), then you can use that relationship if it seems like calculating what you are asked for is complicated or very time-consuming....instead, if it's easier and quicker to find the probability that something will NOT happen, you can try that route and then just subtract from one. It's not that you HAVE to use that approach, but it is generally much faster for questions on the GMAT where time is already an issue.

Note that this is pretty common on questions that ask you to find the probability of "at least"....xyz...when you see that phrase, consider this approach to the problem.
Jen Kedrowski
Kaplan Test Prep & Admissions
GMAT Teacher
MBA Admissions Consultant
National Product Team Member
Join the discussion